# DART: An Automated End-to-End Object Detection Pipeline with Data Diversification, Open-Vocabulary Bounding Box Annotation, Pseudo-Label Review, and Model Training

### v3

@@ 4.5.4 Analysis of YOLO
Finally, the statistics for all real-time object detectors in this study are listed in Table 6. Each YOLO variant is enhanced by the best options provided by each DART component, i.e. photorealistic generated images judged by InternVL-1.5 and mixed with original data at the optimal ratio of 3:1, bounding box annotations carried out by Grounding DINO base model, pseudo-labels approved by the LMMs, and the best hyperparameter set sifting through extensive fine-tuning experiments. The AP values are always derived from the same test set. As seen from the third column, the YOLOv10 models are designed to possess fewer parameters than the YOLOv8 model of the same category (indicated by the suffix n or s). Due to efficiency-driven model design, the YOLOv10n model achieves an AP of 0.802 in just 1.84ms (on a T4 GPU with TensorRT FP16). This is emphasized in the left subfigure of Figure 10, where other models fall slightly behind YOLOv10n in the AP-speed tradeoff. Increasing the number of parameters by approximately 40% and inference time by 30%, the YOLOv8n model significantly boosts the AP to 0.832. The right diagram of Figure 10 clearly shows that YOLOv8n achieves a Pareto optimal balance between AP and model size, outperforming all other models in this regard. If the sole focus is on AP, YOLOv8s becomes the optimal choice, albeit at the cost of the highest parameter count and the slowest inference time. We reasonably speculate that larger and more complex models, such as the m/l/x variants of the YOLO series or state-of-the-art transformer-based object detectors [70, 71] would further elevate AP. However, this contradicts DART’s goal of real-time performance on edge devices, so they are not included in this study.
@@
In summary, the fastest model is YOLOv10n, the most accurate is YOLOv8s, but the overall best model for real-time object detection purposes is YOLOv8n, as it achieves the Pareto optimal balance between AP and model size, with a practical AP performance of 0.832. We propose the trained YOLOv8n as the final output of the DART pipeline. In contrast to the disastrous performance without DARG and the subpar results of an incomplete DART, the comprehensive DART pipeline clearly demonstrates superior efficacy in enhancing the capabilities of real-time object detection models. Compared to the catastrophic performance without DARG and the subpar results of an incomplete DART, the comprehensive DART pipeline clearly demonstrates superior efficacy in enhancing the capabilities of real-time object detection models. The fact that DART accomplishes these advancements without needing manual annotation and extra data collection further highlights its significance in effectiveness and efficiency.
@@ 41. Multiple
Derived from the responses from GPT-4o-based pseudo-label review, Grounding DINO delivers feasible results in even 60% of hard cases with confidence scores below 0.5. Nonetheless, current state-of-the-art open-vocabulary detection (OVD) models still exhibit certain issues. As shown in Figure 12, Grounding DINO’s performance significantly declines when the prompt is simply too long. This limitation prevents us from using fully concatenated class names as a single text prompt (which is the case for the first row in Figure 12 and marked as ”all prompt”). Instead, we must adopt a workaround involving original class names, synonyms, and co-occurring classes and their combinations as detailed in Section 3.3. Additionally, text ambiguity can pose problems. For instance, “handler” (case d in Figure 12) can mean either an object or a person, and “gantry crane” (e) can refer to either the entire structure or individual cranes. Furthermore, OVD is certainly not immune to common challenges faced by all object detection models, including omitted target objects (f), difficulties with small objects (g), incorrect classification (h), or box coordinates(i). Fortunately, LMM can assist in identifying these errors. Our GPT-4o-based pseudo-label review detects 5 out of 6 instances of errors in the last two rows of Figure 12, demonstrating both the necessity and efficacy of the proposed LMM-based review.
Finally, we first visually compare object detection results with and without DART in Figure 13. The first column in the figure visualizes ground truth (labeled by Grounding DINO and approved by GPT-4o), while the other two showcase prediction results of a YOLOv8n model trained with and without DART. Each row picks an image from the test set, which is never seen by the object detector in either scenario. Clearly, without DART’s customized diversification and annotation, the model performance remains poor even after fine-tuning on LVIS. On the contrary, introducing DART leads to a substantial performance boost, yielding almost identical results to the ground truth. This aligns with the dramatic AP increase from 0.064 to 0.832 discussed in Section 4.5.4, demonstrating the necessity and effectiveness of DART.
@@ A.2 LMM
Besides GPT-4o, we have also tried GPT-4v with the ”gpt-4-vision-preview” API and observed that GPT-4v is particularly strict when handling bounding boxes, which leads to the ”relaxation” and ”reminder” section of the user prompt listed in LABEL:tab:prompt_gpt. Additionally, we find that if the prompt is too lengthy, all LMM, except for GPT-4O, experience a decline in performance. Therefore, we deliberately shorten the prompt for InternVL-1.5, as shown in Table 12. For GPT-4v, we discover a workaround for this issue: using multi-turn dialogues, where we first ask the questions (which corresponds to ”task,” ”question,” and ”output” section in LABEL:tab:prompt_gpt) and then give suggestions (”relaxation” and ”reminder” section in LABEL:tab:prompt_gpt)) in the second turn. This approach helps avoid overly long prompts and improves model performance. In contrast, GPT-4o doesn’t suffer from long prompts and dominates its predecessor in the bounding box review task. Therefore, we exclusively use it for the pseudo-label review segment of our DART pipeline.
@@ A.3 YOLO
Table 8 presents hyperparameters used for YOLO models. We follow common practice [17, 20, 18, 19, 21, 22, 23] to resize all images to 640$\times$640. In almost all experiments, training is conducted over 60 epochs since model performance plateaus after 20-30 epochs. AdamW optimizer is always employed with default beta settings and a weight decay set to 5$\times$10-4. For each experiment setup and model type, we meticulously fine-tune learning-rate-related parameters and only display the optimal hyperparameter sets derived from the experiment setup of the optimal generated-to-original data ratio (3:1) in Table 8. It turns out that the optimal learning rates range from 7e-5 to 5e-4, with a general trend indicating that larger models and bigger datasets benefit from lower learning rates. The final learning rate factor is always 50% at best. The learning rate schedule is either cosine or linear, depending on the model. Warm-up epochs are set between 1 and 4, with a warm-up momentum of 0.8 and a warm-up bias learning rate of 0.1. The box loss weight is set to 7.5, class loss weight to 0.5, and DFL loss weight to 1.5. HSV augmentations are specified by saturation with 0.7, value with 0.4, and hue with 0.015. Mosaic augmentation is always applied (with probability 100%) except for the last 10 epochs when it’s turned off for better alignment. Additional parameters include translation at 0.1 fraction of total width and height, horizontal flip rate at 0.5, and scale at 0.5 (resizing between 50% and 150%).
@@ B.1 DreamBooth
LABEL:tab:prompts_dreambooth displays all the prompts used for inference in DreamBooth with SDXL to diversify our dataset. For each prompt, we provide an overall description, the exact prompt text, and their respective IDs. prompts 1-48 are general prompts applied across all classes. Prompts 49-57 are specific to large land-based machinery, such as mining equipment, while prompts 58-67 are tailored for water-related machines, such as maritime cranes.
@@ B.3 LMM
GPT-4o receives detailed prompts for bounding box review. The prompts are broken into parts in LABEL:tab:prompt_gpt, with the purposes given on the left and the exact text contents given on the right. First, The system prompt briefly summarizes the bounding box review task and highlights three key metrics: precision, recall, and fit. The user prompt consists of the five remaining sections. The ”task” section outlines the task and objectives in more detail, with {target} referring to the pre-defined class during data collection and {secondary_target} indicating commonly co-occurring classes also recorded during data collection. Note that explicit mention of target types is not mandatory since the boxes in the images already have class labels, and GPT-4o exhibits excellent OCR capability. But it’s still beneficial for corner cases where the label in the image is occluded. Next, the ”questions” section elaborates on the specific questions corresponding to the three main criteria. The output format is specified in the ”output” segment to facilitate subsequent processing. The ”output” section also follows the guidelines of COT to enhance the model’s performance. In our early experiments, we found that GPT-4 models tend to be overly strict in their judgments regarding the bounding box. To mitigate this, we summarize common issues in the ”relaxation” and ”reminder” sections and attach them to the end of the user prompt. Additionally, our tests with GPT-4v show that including the ”relaxation” and ”reminder” sections as part of the second round of dialogue instead of a single long text is more effective. However, GPT-4o can handle all sections of the user prompt at once.
@@
task In this image, the target object for bounding box annotation is {target}. There may also be {secondary_target}. Each of existing {target}s and {secondary_target}s should be annotated by a bounding box drawn as a colored rectangle. The goal is to accurately localize all {target}s and {secondary_target}s using these bounding boxes. Your task is to evaluate whether all bounding boxes are correct.

### v4

@@ 4.5.4 Analysis of YOLO
Finally, the statistics for all real-time object detectors in this study are listed in Table 6. Each YOLO variant is enhanced using the optimal configurations derived from the DART pipeline. This includes photorealistic synthetic images generated by fine-tuned SDXL models via DreamBooth, subsequently filtered by InternVL-1.5 and mixed with original data at the optimal ratio of 3:1. Bounding box annotations are provided by the Grounding DINO base model, while its output pseudo-labels are validated through GPT-4o. Final hyperparameters are selected via extensive fine-tuning experiments to ensure peak performance. The AP values are always derived from the same test set.
As seen from the third column of Table 6, the YOLOv10 models are designed to possess fewer parameters than the YOLOv8 model of the same category (indicated by the suffix n or s). Due to efficiency-driven model design, the YOLOv10n model achieves an AP of 0.802 in just 1.84ms (on a T4 GPU with TensorRT FP16). This is emphasized in the left subfigure of Figure 10, where other models fall slightly behind YOLOv10n in the AP-speed tradeoff. Increasing the number of parameters by approximately 40% and inference time by 30%, the YOLOv8n model significantly boosts the AP to 0.832. The right diagram of Figure 10 clearly shows that YOLOv8n achieves a Pareto optimal balance between AP and model size, outperforming all other models in this regard. If the sole focus is on AP, YOLOv8s becomes the optimal choice, albeit at the cost of the highest parameter count and the slowest inference time. We reasonably speculate that larger and more complex models, such as the m/l/x variants of the YOLO series or state-of-the-art transformer-based object detectors [70, 71] would further elevate AP. However, this contradicts DART’s goal of real-time performance on edge devices, so they are not included in this study.
@@
In summary, the fastest model in our experiments is YOLOv10n, the most accurate one is YOLOv8s, but the overall best model for real-time object detection purposes is YOLOv8n, as it achieves the Pareto optimal balance between AP and model size, with a practical AP performance of 0.832. We propose the trained YOLOv8n as the final component of the DART pipeline. In contrast to the disastrous performance without DARG and the subpar results of an incomplete DART, the comprehensive DART pipeline clearly demonstrates superior efficacy in enhancing the capabilities of real-time object detection models. The fact that DART accomplishes these advancements without manual annotation and extra data collection further highlights its significance in effectiveness and efficiency.
@@ 41. Multiple
Derived from the responses from GPT-4o-based pseudo-label review, Grounding DINO delivers feasible results in even 60% of hard cases with confidence scores below 0.5. Nonetheless, current state-of-the-art open-vocabulary detection (OVD) models still exhibit certain issues. As shown in Figure 12, Grounding DINO’s performance significantly declines when the prompt is simply too long. This limitation prevents us from using fully concatenated class names as a single text prompt (which is the case for the first row in Figure 12 and marked as ”all prompt”). Instead, we must adopt a workaround involving original class names, synonyms, and co-occurring classes and their combinations as detailed in Section 3.3. Additionally, text ambiguity can pose problems. For instance, “handler” (case d in Figure 12) can mean either an object or a person, and “gantry crane” (e) can refer to either the entire structure or individual cranes. Furthermore, OVD is certainly not immune to common challenges faced by all object detection models, including omitted target objects (f), difficulties with small objects (g), incorrect classification (h), and bounding box coordinates (i). Fortunately, LMM can assist in identifying these errors. Our GPT-4o-based pseudo-label review detects 5 out of 6 instances of errors in the last two rows of Figure 12, demonstrating both the necessity and efficacy of the proposed LMM-based review.
Finally, we visually compare object detection results with and without DART in Figure 13. The first column in the figure visualizes ground truth (labeled by Grounding DINO and approved by GPT-4o), while the other two showcase prediction results of a YOLOv8n model trained with and without DART. Each row picks an image from the test set, which is never seen by the object detector in either scenario. Clearly, without DART’s customized diversification and annotation, the model performance remains poor even after fine-tuning on LVIS. On the contrary, introducing DART leads to a substantial performance boost, yielding almost identical results to the ground truth. This aligns with the dramatic AP increase from 0.064 to 0.832 discussed in Section 4.5.4, demonstrating the necessity and effectiveness of DART.
@@ A.2 LMM
Besides GPT-4o, we have also tried GPT-4v with the ”gpt-4-vision-preview” API and observed that GPT-4v is particularly strict when handling bounding boxes, which leads to the ”relaxation” and ”reminder” section of the user prompt listed in LABEL:tab:prompt_gpt. Additionally, we find that if the prompt is too lengthy, all LMMs, except for GPT-4o, experience a decline in performance. Therefore, we deliberately shorten the prompt for InternVL-1.5, as shown in Table 12. For GPT-4v, we discover a workaround for this issue: using multi-turn dialogues, where we first ask the questions (which corresponds to ”task,” ”question,” and ”output” section in LABEL:tab:prompt_gpt) and then give suggestions (”relaxation” and ”reminder” section in LABEL:tab:prompt_gpt)) in the second turn. This approach helps avoid overly long prompts and improves model performance. In contrast, GPT-4o doesn’t suffer from long prompts and dominates its predecessor in the bounding box review task. Therefore, we exclusively use it for the pseudo-label review segment of our DART pipeline.
@@ A.3 YOLO
Table 8 presents hyperparameters used for YOLO models. We follow common practice [17, 20, 18, 19, 21, 22, 23] to resize all images to 640$\times$640. In almost all experiments, training is conducted over 60 epochs since model performance plateaus after 20-30 epochs. AdamW optimizer is always employed with default beta settings and a weight decay set to 5$\times$10-4. For each experiment setup and model type, we meticulously fine-tune learning-rate-related parameters and only display the optimal hyperparameter sets derived from the experiment setup of the optimal generated-to-original data ratio (3:1) in Table 8. It turns out that the optimal learning rates range from 7e-5 to 5e-4, with a general trend indicating that larger models and bigger datasets benefit from lower learning rates. The final learning rate factor is always 50% at best. The learning rate schedule is either cosine or linear, depending on the model. Warm-up epochs are set between 1 and 4, with a warm-up momentum of 0.8 and a warm-up bias learning rate of 0.1. The box loss weight is set to 7.5, the class loss weight to 0.5, and the DFL loss weight to 1.5. HSV augmentations are specified by saturation with 0.7, value with 0.4, and hue with 0.015. Mosaic augmentation is always applied (with probability 100%) except for the last 10 epochs when it’s turned off for better alignment. Additional parameters include translation at 0.1 fraction of total width and height, horizontal flip rate at 0.5, and scale at 0.5 (resizing between 50% and 150%).
@@ B.1 DreamBooth
LABEL:tab:prompts_dreambooth displays all the prompts used for inference in DreamBooth with SDXL to diversify our dataset. For each prompt, we provide an overall description, the exact prompt text, and its respective ID. Prompts 1-48 are general prompts applied across all classes. Prompts 49-57 are specific to large land-based machinery, such as mining equipment, while prompts 58-67 are tailored for water-related machines, such as maritime cranes.
@@ B.3 LMM
GPT-4o receives detailed prompts for bounding box review. The prompts are broken into parts in LABEL:tab:prompt_gpt, with the purposes given on the left and the exact text contents on the right. First, the system prompt briefly summarizes the bounding box review task and highlights three key metrics: precision, recall, and fit. The user prompt consists of the five remaining sections. The ”task” section outlines the task and objectives in more detail, with {target} referring to the pre-defined class during data collection and {secondary_target} indicating commonly co-occurring classes also recorded during data collection. Note that explicit mention of target types is not mandatory since the boxes in the images already have class labels, and GPT-4o exhibits excellent OCR capability. But it’s still beneficial for corner cases where the label in the image is occluded. Next, the ”questions” section elaborates on the specific questions corresponding to the three main criteria. The output format is specified in the ”output” segment to facilitate subsequent processing. The ”output” section also follows the guidelines of COT to enhance the model’s performance. In our early experiments, we found that GPT-4 models tend to be overly strict in their judgments regarding the bounding box. To mitigate this, we summarize common issues in the ”relaxation” and ”reminder” sections and attach them to the end of the user prompt. Additionally, our tests with GPT-4v show that including the ”relaxation” and ”reminder” sections as part of the second round of dialogue instead of a single long text is more effective. However, GPT-4o can handle all sections of the user prompt at once.
@@
task In this image, the target object for bounding box annotation is {target}. There may also be {secondary_target}. Each of the existing {target}s and {secondary_target}s should be annotated by a bounding box drawn as a colored rectangle. The goal is to accurately localize all {target}s and {secondary_target}s using these bounding boxes. Your task is to evaluate whether all bounding boxes are correct.